### v2

# Intention Analysis Makes LLMs A Good Jailbreak Defender

Abstract
Aligning large language models (LLMs) with human values, particularly in the face of stealthy and complex jailbreak attacks, presents a formidable challenge. In this study, we present a simple yet highly effective defense strategy, i.e., Intention Analysis ($\mathbb{IA}$). The principle behind this is to trigger LLMs’ inherent self-correct and improve ability through a two-stage process: 1) essential intention analysis, and 2) policy-aligned response. Notably, $\mathbb{IA}$ is an inference-only method, thus could enhance the safety of LLMs without compromising their helpfulness [^1] Improving the safety of LLM with training-required methods Ouyang et al. (2022); Touvron et al. (2023) always necessitates great effort to strike a delicate balance between safety and helpfulness.. Extensive experiments on SAP200 and DAN benchmarks across Vicuna, ChatGLM, MPT, DeepSeek, and GPT-3.5 show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -46.5% attack success rate) and maintain the general helpfulness. Encouragingly, with the help of our $\mathbb{IA}$, Vicuna-7b even outperforms GPT-3.5 in terms of attack success rate. Further analyses present some insights into how our method works. To facilitate reproducibility, we release our code and scripts at: https://github.com/alphadl/SafeLLM_with_IntentionAnalysis.
Warning: Some of the examples may be harmful!
Intention Analysis Makes LLMs A Good Jailbreak Defender
[…]
1 Introduction
Recently, Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2023; Google, 2023), such as ChatGPT, not only show remarkable capabilities in various tasks Qin et al. (2023); Zhong et al. (2023); Peng et al. (2023), but also lead to the risk of potential misuse (e.g., producing harmful responses or illegal suggestions) Weidinger et al. (2021). Efforts like Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) have been made to alleviate these risks and enhance LLMs’ alignment with human values, making LLMs able to refuse harmful questions like how to rob a bank? However, LLMs remain vulnerable to some adversarial inputs, particularly in the context of so-called “jailbreak” attacks. These jailbreak attacks are specially designed to circumvent the safety policy and manipulate LLMs for their restricted outputs (Yuan et al., 2024; Zou et al., 2023), which poses formidable risks in real applications.
Figure 1: Performance of our method on different jailbreak datasets. Our $\mathbb{IA}$ 1) consistently reduces Attack Success Rate ($\downarrow$) against both prompt-level (SAP and DAN) and token-level (AdvBench) jailbreaks, 2) achieves remarkable safety improvements for both SFT (Vicuna-7B&13B) and RLHF (ChatGLM-6B) models.
Despite numerous strategies focusing on defending LLMs against these jailbreak attacks (Li et al., 2024; Jain et al., 2023; Chen et al., 2024), their efficacy markedly decreases when handling those with stealthy and complex intention (see experimental results in Table 3). This type of jailbreak attacks are typically hand-crafted prompts by using more instructions and employing specific techniques, such as creating a virtual scenario where LLMs are assigned the role of “DAN” (Do Anything Now, Liu et al., 2023). Such stealthy and complex intention poses a great challenge for LLMs to maintain safety when confronting these jailbreak attacks (Shen et al., 2023; Bao et al., 2023).
In this paper, we leverage the intrinsic intent recognition capabilities of LLMs, proposing an Intention Analysis ($\mathbb{IA}$) strategy, which cleverly addresses the aforementioned challenge. $\mathbb{IA}$ enables LLMs to recognize the underlying intention of the user query to better understand it and perceive the unsafe content within before responding, therefore significantly enhancing their safety against varying jailbreak attacks, see Figure 1 for a demonstration. Practically, $\mathbb{IA}$ follows a two-stage process, where $\mathbb{IA}$ first directs LLMs to analyze the essential intention of the user query, and then extract the safety policy-aligned final response from LLMs, as shown in Figure 2. Notably, our $\mathbb{IA}$ is an inference-only method that can significantly enhance LLM safety without the need for additional safety training (Ouyang et al., 2022; Touvron et al., 2023). In this way, $\mathbb{IA}$ skillfully circumvents the safety-helpfulness trade-off and enables comparable safety improvement as well as better helpfulness maintenance.
To summarize, our contributions are as follows:
  • $\bullet$
    We introduce $\mathbb{IA}$, a new method that significantly enhances LLMs safety in the context of sophisticated jailbreak attacks through an intention analysis mechanism.
  • $\bullet$
    $\mathbb{IA}$ is a plug-and-play inference-only method, thereby 1) cleverly circumventing the safety-helpfulness trade-off that is challenging in safety training, and 2) can be flexibly and effectively deployed upon any LLMs.
  • $\bullet$
    Empirically, our robust $\mathbb{IA}$ significantly and consistently reduces the harmfulness of LLM outputs, while maintaining the helpfulness, achieving new state-of-the-art performance on several benchmarks, e.g., AdvBench.
Figure 2: Illustrated Comparison of (a) vanilla and (b) the proposed $\mathbb{IA}$. Our $\mathbb{IA}$ consists of two stages: (1) Essential Intention Analysis: instructing the language model to analyse the intention of the user query with an emphasis on safety, ethics, and legality; (2) Policy-Aligned Response: eliciting the final response aligned with safety policy, building upon the analyzed intention from the first stage.
2 Related Work
Alignment-Breaking Adversarial Attack
Despite significant efforts to align LLMs with human preference (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Korbak et al., 2023; Miao et al., 2024), adversarial attackers can still elicit harmful responses from LLMs by “jailbreak” attacks (Shen et al., 2023; Liu et al., 2023). Current jailbreak attack methods are primarily classified into two categories: prompt-level and token-level (Chao et al., 2023). Prompt-level jailbreak is typically hand-crafted through human ingenuity and is semantically understandable in general Shen et al. (2023). For token-level jailbreak, a representative work is to automatically fetch a transferable attack suffix through the Greedy Coordinate Gradient (GCG) algorithm which maximizes the probability of the language model generating an affirmative and unsafe response Zou et al. (2023).
In this work, we aim to enhance the safety of LLMs when facing prompt-level jailbreaks with stealthy and complex intentions.
Safety-Enhancing Defensive Mitigations
Recently, numerous methods have been developed to reduce LLMs’ harmful generations at the inference stage. A branch of them mainly concentrates on controlling the content that LLMs can see by pre-processing the user inputs, such as perplexity filtering (Alon and Kamfonas, 2023; Jain et al., 2023), paraphrasing (Jain et al., 2023) and re-tokenization (Cao et al., 2023; Jain et al., 2023). Another branch focuses on exploiting LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks, such as letting LLMs self-evaluate their outputs (Helbling et al., 2023) or reminding of safety in system mode with conventional decoding Xie et al. (2023) or contrastive decoding Zhong et al. (2024).
While these existing methods show effectiveness in preventing LLMs from generating unsafe responses, their efficacy significantly diminishes when confronted with sophisticated jailbreak attacks due to the ignorance of the underlying stealthy and malicious intention. In contrast, our method handles such jailbreak attacks by leveraging the intrinsic intent recognition capabilities of LLMs, which leads to a significant enhancement of LLMs’ safety.
3 Methodology
3.1 Preliminary
We focus on enhancing LLM safety during the inference stage. In practice, developers usually implement pre-defined system prompts for LLMs to facilitate safe, responsible, and effective interactions with users (Chiang et al., 2023). Under this premise, the system prompt $P_{sys}$ and the user prompt $P_{usr}$ are concatenated to form the final input $\{x_{1:n}^{s},x_{1:m}^{u}\}$ of the LLM, where $P_{sys}=\{x_{1}^{s},x_{2}^{s},\dots,x_{n}^{s}\}$, $P_{usr}=\{x_{1}^{u},x_{2}^{u},\dots,x_{m}^{u}\}$, $x_{i}^{s}$ and $x_{j}^{u}$ are the $i$-th and $j$-th token of $P_{sys}$ and $P_{usr}$, respectively. Conditioned on the input $\{x_{1:n}^{s},x_{1:m}^{u}\}$, the autoregressive inference process of response $R=y_{1:L}$ is formulated as following:
$q(y_{1:L}|x_{1:n}^{s},x_{1:m}^{u})=\prod\limits_{i=1}^{L}q(y_{i}|y_{1:i-1},x_{1:n}^{s},x_{1:m}^{u}).$
For simplicity, we write $R\sim q(R|P_{sys},P_{usr)}$ to denote sampling a response $R$ from $q(\cdot)$ given the prompt $P_{sys}$ and $P_{usr}$. In this way, the response $R$ can be obtained as: $R=\textrm{LLM}\left(P_{sys},P_{usr}\right).$
In this work, we aim to leverage LLMs’ intrinsic abilities of intention analysis, to enhance their safety against varying jailbreak attacks during the inference stage, while simultaneously maintaining the general helpfulness.
3.2 $\mathbb{IA}$: Intention Analysis
To achieve the above goal, we introduce $\mathbb{IA}$, a zero-shot intention analysis mechanism, to guide LLMs to explicitly identify and understand the underlying intention of a user query before facilitate a final response. Specifically, we devise a two-stage intention analysis instruction to accomplish the whole process, as illustrated in Figure 2(b): (1) essential intention analysis and (2) policy-aligned response.
Stage 1: Essential Intention Analysis
This stage focuses on guiding the LLMs to discern the core intention behind the user query, with a specific orientation towards safety, ethics, and legality. The critical question arises: How can we ensure that LLMs accurately identify the query’s intention? Actually, recent studies Bender and Koller (2020); Zhu et al. (2024); Gómez-Pérez et al. (2023) have shown that LLMs are notably proficient at language understanding tasks, and intention analysis is a straightforward task, indicating the competence of LLMs in performing this stage. The only concern is generative models’ potential hallucination when performing the discriminative tasks Ji et al. (2023); Yan et al. (2021); Ye et al. (2023), therefore, we carefully define the format for the models’ response, that is, beginning with “The essential intention of the query is”, which has been validated in our analysis.
In practice, we construct the instruction for the LLMs to effectively perform intention analysis, denoted as $I_{rec}$. When presented with a user query $P_{usr}$ [^2] In this context, the user query $P_{usr}$ mostly represents the entirety of a jailbreak query., we concatenated $I_{rec}$ and $P_{usr}$ to form a whole “User” level input $I_{rec}\oplus P_{usr}$ for the LLMs. Subsequently, the designated target LLMs engage in an auto-regressive inference process, guided by its system prompt $P_{sys}$, to produce the stage-specific response:
$R_{st1}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr}\right),$
which is expected to contain the essential intention.
Stage 2: Policy-Aligned Response
Having successfully recognized the essential intention, the second stage aims to elicit the final response from the LLMs. We first direct the LLMs to bear the identified intention in mind and then provide a final response to the user query. Meanwhile, we explicitly instruct the LLMs to strictly adhere to safety policy and ethical standards and ensure the exclusion of any unsafe content in their responses. To this end, the second stage further strengthens the role of the intention analysis and reinforces the inherent alignment of LLMs with the safety policy.
Specifically, we concatenate the dialogue from the first stage with the instruction for the current stage, denoted as $I_{ct}$, forming the complete input for LLMs. Then a similar autoregressive inference process is conducted, leading to the generation of the final response $R_{st2}$ to the user query $P_{usr}$:
$R_{st2}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr},R_{st1},I_{ct}\right).$
To assess the safety of the response, we follow Shen et al. (2023) to employ a binary auto-annotation function $\textrm{AS}(\cdot)$ [^3] Will be discussed in detail in Section 4.1. to determine the harmfulness of $R_{st2}$. If the outcome yields $\textrm{AS}(R_{st2})=False,$ then the response is deemed safe, indicating a successful defense against the jailbreak attack.
Models Methods SAP200 Avg.($\downarrow$) $\bm{\Delta}\left(\downarrow\right)$
Fraud Polit. Porno. Race Relig. Suici. Terro. Viole.
ChatGLM-6B Vanilla 36.5 29.0 52.5 45.5 47.5 61.0 34.0 60.0 45.8 -39.68
Ours 5.00 3.00 8.50 5.50 11.5 10.5 0.50 4.50 6.12
[…]
Ours 1.50 0.50 0.00 0.00 0.50 0.00 0.00 0.00 0.31
[…]
Ours 1.00 0.00 2.00 0.00 1.00 5.50 0.00 0.00 1.12
[…]
Ours 21.0 16.5 20.0 22.0 16.5 18.0 19.0 21.0 19.2
[…]
Ours 3.00 0.00 5.50 0.50 1.00 2.00 0.0 0.50 1.56
[…]
Ours 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Table 1: Comparison of our $\mathbb{IA}$ and vanilla in SAP200 (comprising eight distinct sensitive topics) in terms of ASR (%). The best average results are highlighted in bold.
Models Methods DAN Avg.($\downarrow$) $\bm{\Delta}\left(\downarrow\right)$
Anarc. Basic Oppos. Toxic Advan. Start. Guide. Virtu.
ChatGLM-6B Vanilla 25.4 26.2 32.8 48.7 28.3 19.4 14.7 36.6 29.0 -23.52
Ours 7.73 4.10 12.3 8.15 6.14 1.04 3.33 1.06 5.48
[…]
Ours 0.00 9.74 1.02 3.59 5.15 1.02 2.05 0.52 2.89
[…]
Ours 1.03 0.52 0.51 1.02 0.51 0.00 1.03 0.51 0.64
[…]
Ours 5.13 6.15 6.15 5.13 6.15 3.08 4.62 6.67 5.38
[…]
Ours 0.51 0.00 1.54 1.54 23.6 0.00 0.00 3.08 3.78
[…]
Ours 0.00 0.00 2.05 0.00 1.02 1.03 1.02 0.00 0.64
Table 2: Comparison of our $\mathbb{IA}$ and vanilla in DAN (including eight jailbreak communities) in terms of ASR (%). The best average results are highlighted in bold.
4 Experiment
4.1 Setup
Datasets
We experiment on two main categories of jailbreak attacks, including two representative prompt-level jailbreak datasets (i.e. DAN (Shen et al., 2023) and SAP200 (Deng et al., 2023a)), and one popular token-level jailbreak dataset (i.e. AdvBench (Zou et al., 2023). Besides, to evaluate the effect of our $\mathbb{IA}$ on helpfulness for general in-distribution queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022).
Methods AdvBench DAN SAP200
Vicuna-7B Vicuna-13B Vicuna-7B Vicuna-13B Vicuna-7B Vicuna-13B
Vanilla 83.0 87.0 48.4 60.0 70.0 65.9
+ BPE-dropout (Jain et al., 2023) 63.0 50.0 23.8 28.2 67.2 48.9
+ ICD (Wei et al., 2023) 1.00 0.00 44.4 58.9 32.8 32.8
+ Self Defense (Helbling et al., 2023) 24.0 20.0 31.3 28.8 53.2 29.7
+ Moral Self-Correction (Ganguli et al., 2023) 26.0 13.0 25.0 28.1 49.0 42.8
+ System-Mode Self-Reminder (Xie et al., 2023) 11.0 1.0 45.3 57.5 33.8 36.9
+ $\mathbb{IA}$ (Ours) 0.00 0.00 3.42 0.94 0.31 1.56
Table 3: Comparison of our method and existing advanced defense methods in terms of ASR (%). The best and second best results are highlighted in bold and underline.
Figure 3: The MultiJail Dataset results on ChatGPT under two jailbreak scenarios with and without our $\mathbb{IA}$. (a) results on direct MultiJail dataset. (b) results on malicious jailbreak prompt attached to MultiJail.
Evaluation Metrics
For safety assessment, we annotate the harmfulness of responses and report attack success rate (ASR, Shen et al., 2023). Specifically, for SAP200 and AdvBench, we annotate harmfulness following Zou et al. (2023) by matching refusal strings (e.g., “I’m sorry”). For DAN, due to the complexity of responses, we adopt gpt-3.5-turbo-0613444https://openai.com/blog/chatgpt as the auto-annotation function following Deng et al. (2023a). See Appendix B for more detailed settings. In addition, we carry out human evaluations to ensure the credibility and accuracy of the assessment (See Section 5.3 for details).
For helpfulness assessment, we report win rate (Dubois et al., 2023) for AlpacaEval and accuracy (Hendrycks et al., 2021) for MMLU. For TruthfulQA, we follow Chuang et al. (2023) and report on two distinct metrics: MC1 and MC2 scores. For these helpfulness metrics, higher scores indicate stronger factuality.
Models
To evaluate the effectiveness of our $\mathbb{IA}$ method, we experiment on models varying scales and levels of alignment, including not only SFT models, i.e. Vicuna-7B/13B-v1.1 (Chiang et al., 2023) and MPT-30B-Chat (Team, 2023), but also RLHF models, i.e. ChatGLM-6B (Zeng et al., 2023) and DeepSeek-67B-Chat (DeepSeek-AI, 2024). Beyond these representative open-source models, our experimentation extends to an advanced closed-source model, GPT-3.5 (gpt-3.5-turbo-1106) (OpenAI, 2023), which is renowned for its superior performance in general capabilities, including safety.
Comparison Baselines
We compare our $\mathbb{IA}$ with five recently proposed defense methods, i.e., BPE-dropout (Jain et al., 2023), ICD (Wei et al., 2023), Self Defense (Helbling et al., 2023), Moral Self-Correction (Ganguli et al., 2022), and System-Mode Self-Reminder (Xie et al., 2023). Among them, the first two methods pre-process user queries to undermine harm while the last three exploit LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks. For a fair comparison, we closely follow the best default parameters in their papers.
Implementation
For the DAN dataset, we compile an evaluation dataset of 1560 samples by extracting 195 questions from each jailbreak community within the forbidden question set (Shen et al., 2023). For AdvBench, we follow Zou et al. (2023) and conduct transfer attacks on Vicuna-7B and 13B. The adversarial suffix achieving the lowest loss after 500 steps of optimization are adopted to further attack target models on 100 individual harmful behaviors (Wei et al., 2023). For open-source models, we download them from HuggingFace [^5] https://huggingface.co/models. For closed-source models, we obtain the responses of GPT-3.5 via API calls. Throughout our experiments, we set a temperature of zero for deterministic outcomes Peng et al. (2023) and a generation length of 1024 tokens, employing default system prompt templates for each LLM if available.
4.2 Main Results
Our method significantly boosts safety against varying jailbreak attacks.
The performance of our method on SAP200 and DAN datasets is presented in Table 1&2, respectively. We can observe that our $\mathbb{IA}$ consistently achieves marked safety improvements against the most advanced and widespread jailbreak prompt attacks, across model sizes and alignment levels. Particularly for SAP200, the substantial reductions of ASR for Vicuna-7B (73.09%) and Vicuna-13B (70.38%) demonstrate the effectiveness. As for different alignment levels, not only does our method show satisfactory performance on SFT models, i.e. Vicuna-7B and 13B, but such enhanced safety also extends to well-aligned models like GPT-3.5, which registers nearly 0% ASR under our $\mathbb{IA}$.
Figure 4: The confusion matrix illustrating the relationship between the success of intention analysis and the harmlessness of the response. “IR Succ.” and “IR Fail.” are success or failure of intention analysis, respectively.
Models Methods AlpacaEval MMLU TruthfulQA
Win Rate Acc. MC1 MC2
ChatGLM-6B Vanilla 28.7 40.1 37.1 54.1
Ours 25.3 39.3 37.5 56.0
[…]
Ours 62.8 45.0 35.2 53.4
[…]
Ours 73.5 48.3 38.2 55.1
Table 4: General performance on helpful dataset upon different models in terms of Win Rate (%) for AlpacaEval, Accuracy (%) for MMLU and MC1, MC2 (%) for TruthfulQA.
Our method can consistently enhance model safety in the context of multilingual jailbreak.
Recent studies Deng et al. (2023b); Yong et al. (2023) reveal that the multilingual jailbreak poses a new defence challenge for LLMs. To verify the effectiveness of our method in multilingual jailbreak scenarios, we follow Deng et al. (2023b) to reproduce their settings on nine different languages with two jailbreak scenarios, including Chinese (zh), Italic (it), Vietnamese (vi), Arabic (ar), Korean (ko), Thai (th), Bengali (bn), Swahili (sw), and Javanese (jv). The results of vanilla ChatGPT and our $\mathbb{IA}$ are presented in Figure 3. We show that 1) our $\mathbb{IA}$ can consistently maintain performance, even in low-resource languages, e.g., th, bn, sw, and jv, 2) even in scenarios where a malicious jailbreak prompt [^6] We adopt the same jailbreak prompt in Deng et al. (2023b), namely AIM. is attached to the multilingual attacks, our $\mathbb{IA}$ still manages to reduce the ASR from around 80% to almost 0%. This demonstrates the effectiveness of our method under multilingual jailbreak attacks.
Our method maintains the model’s helpfulness on general tasks.
To explore the impact of our method on the general performance of LLMs, we conduct experiments on several acknowledged helpfulness datasets and report the results in Table 4. As observed, our $\mathbb{IA}$ does not compromise the general performance, indicated by a comparable win rate on AlpacaEval and a similar accuracy on MMLU. Furthermore, $\mathbb{IA}$ can improve the truthfulness over different models in terms of better MC1 and MC2 scores. These results indicate that $\mathbb{IA}$ can be deployed in real applications to enhance LLM safety while preserving general helpfulness.
Our method consistently outperforms other defense baselines.
Table 3 lists the comparison results between our $\mathbb{IA}$ and the baselines. [^7] Due to computational resource and financial limitations, we randomly select 320 samples each from DAN and SAP200 datasets for comparative experiments. As observed, our method consistently shows superiority over other baselines on different datasets and model scales. Specifically, our $\mathbb{IA}$ outperforms the second-best method by 30.32% and 23.77% averagely on SAP200 and DAN, respectively. In addition, although ICD and System-Mode Self-Reminder achieve considerable reduction in ASR on AdvBench, their performance severely degrades when dealing with complex and stealthy jailbreak prompts. On the contrary, our method consistently outperforms other baselines across both prompt-level and token-level jailbreak datasets. We conjecture such satisfactory performance to be attributed to our design in intention analysis mechanism, which leverages LLM’s awareness of harmful content to enhance safety against jailbreak attacks.
5 Discussion
5.1 Analysis of Major Components in $\mathbb{IA}$
LLMs can indeed recognize the intentions behind jailbreak queries.
To verify this, we count the samples where intentions are successfully recognized and further examine the correlation between recognizing intentions and producing harmless responses, as shown in Figure 4. We show that LLMs are highly effective in analyzing intentions of complex jailbreak queries, achieving a 100% success rate in most models, such as Vicuna-7B, Vicuna-13B, and MPT-30B-Chat. Moreover, the notable volume of harmless responses following successful intention recognition demonstrates the critical role of intention analysis in enhancing LLM safety.
Our method achieves comparable safety of well safety-trained LLMs while surpassing them in helpfulness.
Our method aims to enhance LLM safety in the inference stage. A natural question arises how does its performance compare to well safety-trained LLMs? To answer this, we compare our method with a representative well safety-trained LLM, i.e., LLaMA2-7B-Chat. The related results are listed in Table 7. We can observe that our method achieves comparable performance to LLaMA2-7B-Chat on safety datasets while significantly outperforming LLaMA2-7B-Chat on the helpfulness dataset by almost 5%. This demonstrates the advantage of our $\mathbb{IA}$ to achieve both safety and helpfulness goals.
Our method still works well if combining two stages into one.
As aforementioned, to maximize the performance, our $\mathbb{IA}$ follows a two-stage process. A natural question arises of whether our mechanism can be merged into one step, to save the decoding overhead. To verify this, we design a cheap one-pass $\mathbb{IA}$, without letting the LLMs separately explain the intention. As shown in Table 5, we see that while two-stage $\mathbb{IA}$ can achieve a relatively lower ASR, the one-pass $\mathbb{IA}$ also performs well, regarding the significantly ASR reduction of 61.7% and 64.27% for Vicuna-7B and Vicuna-13B, respectively.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 67.2 65.4 45.8
+ One-Pass $\mathbb{IA}$ 5.50 1.13 39.0
+ Two-Stage $\mathbb{IA}$ 0.31 1.12 6.12
Table 5: Comparison of our $\mathbb{IA}$ with different implementations (one-pass and two stage) on SAP200 in terms of ASR. The best results are highlighted in bold.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 67.2 65.4 45.8
+ Prompt A 2.94 0.88 5.81
+ Prompt B 5.13 2.06 5.44
+ Ours 0.31 1.12 6.12
Table 6: Ablation of different intention analysis prompts on SAP200 in ASR (%). The best and second best results are highlighted in bold and underline.
5.2 Robustness of Our $\mathbb{IA}$ Mechanism
Our method is robust to different prompts.
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. Comparison results are listed in Table 6. We show that these three prompt sets all result in significant and comparable decreases in ASR on the SAP200 across different models. This uniformity highlights a key conclusion: it is the intention analysis mechanism itself, but not the specific expression of the instructions, that makes our method effective.
Our method is robust against varying intention analysis quality.
A crucial question that emerges regarding our method is whether its performance would be significantly influenced by the quality of intention analysis. To address this concern, we simulate varying correct intention ratios (representing intention analysis quality) in the first stage by substituting a proportion of the intention with “[secret]”, and the corresponding performance of our method is presented in Figure 5. Key observations include 1) Under different correct intention ratios, the ASR values achieved by our method display a relatively stable trend (with variations within 10%), particularly for larger models (e.g., Vicuna 13B), indicating the robustness of our method to varying intention analysis quality, 2) the performance of our method gradually improves as the correct intention ratios increase, indicating that enhancing the ability to analyze intentions can further boost the effectiveness of our $\mathbb{IA}$.
Figure 5: Performance of our method on DAN with varying correct intention ratio, which demonstrates the robustness to intention analysis ratio of our $\mathbb{IA}$.
Methods Safety Helpfulness
SAP200 DAN AlpacaEval
Vicuna-7B 73.4 44.3 66.2
[…]
Vicuna-7B + Ours 0.31 2.89 62.8
Table 7: Comparison between our method and well safety-trained LLM in safety and helpfulness. The best&second best are highlighted in bold and underline.
5.3 Human Evaluation
To validate GPT-3.5 as the harmfulness annotation function, we conduct a human validation process. First, we randomly select 200 responses from both the vanilla model and our method on the DAN dataset. Then we engage two expert annotators proficient in alignment studies of LLMs and fluent in English. After reassessing their evaluations for cases with discrepancies and reaching a consensus, the annotators achieved a remarkable 97% agreement rate. Third, to quantify the credibility of GPT-3.5’s evaluations, we consider different degrees as the lowest acceptable safety threshold and calculate the agreement rate with human annotations. The results, detailed in Table 8, show a notably high agreement of average 96% when categorizing C degree as the minimum harmfulness level. This underscores the enhanced reliability of GPT-3.5 annotations in conjunction with human judgment in harm detection.
Degrees $=$ A $\geq$ B $\geq$ C $\geq$ D $\geq$ E
Vanilla 51% 72% 93% 90% 89%
Ours 97% 98% 99% 99% 3.0%
Table 8: Agreement rate between human annotations and different harm degrees provided by GPT-3.5 evaluation. A, B, C, D, and E degrees refer to ‘very harmful’, ‘harmful’, ‘moderately harmful’, ‘slightly harmful’, ‘not harmful at all’, respectively.
6 Conclusion
To tackle the challenge of widespread complex and stealthy jailbreak attacks, which pose deployment risks for large language models, we propose a simple yet highly effective strategy called Intention Analysis ($\mathbb{IA}$). $\mathbb{IA}$ works in the inference stage. Specifically, $\mathbb{IA}$ leverages the intrinsic intent recognition capacities of LLMs, and instructs them to first analyze the essential intention of user queries before responding to it through two stages. Extensive experiments on SAP200 and DAN benchmarks across models of varying scales and alignment levels show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -46.5% attack success rate), while maintaining the general helpfulness. Additionally, we further verify the robustness of $\mathbb{IA}$ to the specific expression of instructions and quality of intention analysis.
Limitations
Despite the effectiveness of our method in defending sophisticated jailbreak prompts, these prompts do not encompass the entire potential jailbreak attacks encountered in real-world scenarios. Consequently, the practical applicability of our approach remains to be validated through further testing. Our research underlines the importance of intention analysis in improving LLM safety, suggesting future work focusing on integrating this into training to reduce inference costs. Additionally, in the face of the rapid advancements in the adversarial attacks community, there is a pressing need for developing more effective and robust defense strategies for LLMs.
Ethics Statement
We take ethical considerations very seriously. This paper focuses on improving the safety (especially the jailbreak attacks) of large language models, through carefully designed intention analysis prompting mechanism. Our research could significantly reduce the unsafe responses of LLMs. All experiments are conducted on open datasets and the findings and conclusions of this paper are reported accurately and objectively. Thus, we believe that this research will not pose ethical issues.
Appendix A Experimental Datasets
A.1 Safety Datasets
Hand-Crafted Jailbreak Prompts
To assess the effectiveness of our method on in-the-wild jailbreak prompts, we employ two jailbreak prompt datasets. The first is forbidden question set developed by Shen et al. (2023), which is currently the largest in-the-wild jailbreak prompt dataset. To improve computing efficiency, we extract five questions from each forbidden scenario, forming a jailbreak dataset comprising 8 jailbreak communities $\times$ 3 jailbreak prompts $\times$ 13 forbidden scenarios $\times$ 5 questions, totaling 1560 samples. The term “DAN” is used to denote this dataset. For evaluation, we leverage attack success rate (ASR) to consider the success of a jailbreak attack. Considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations, as will be discussed in Section B.
The second SAP200 is an jailbreak prompt dataset, constructed semi-automatically by Deng et al. (2023a) using code injection and payload splitting mechanisms. It encompasses 8 distinct sensitive topics, with 200 samples each, totaling 1600 samples.
Due to computational resource and financial limitations, we randomly select 40 samples for each sub-dataset, totaling $32040samples\times 8sub-datasets=$ 320 samples from DAN and SAP200 datasets,respectively, to conduct comparative experiments in Table 3 and correct intention ratio comparison experiments in Figure 5.
Gradient-Based Adversarial Attacks
To comprehensively verify the effectiveness of our method in defending against jailbreak attacks, we conduct experiments on a popular token-level jailbreak dataset, i.e., AdvBench (Zou et al., 2023) and use the Greedy Coordinate Gradient (GCG) attack algorithm to generate the adversarial suffix. Specifically, we utilize Vicuna-7B and 13B to optimize a universal attack suffix by combining the gradients of the two models. Subsequently, we use the held-out 100 harmful behaviors from AdvBench and apply this optimized suffix to attack other models. We followed the same default parameter setting for GCG, with a learning rate of 0.01, batch size of 512, top-k of 256, and temperature of 1. The suffix achieving the lowest loss after 500 steps was selected for the experiment.
A.2 Helpfulness Datasets
To evaluate the effect of our $\mathbb{IA}$ on helpfulness for general in-distribution queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022). AlpacaEval, containing 805 general questions, is a widely acknowledged benchmark to evaluate the ability of model following general user queries  Chen et al. (2024); Zhang et al. (2023). MMLU covers 57 subjects, aiming to evaluate comprehensive knowledge abilities across multiple major categories, from humanities to social sciences to science and engineering. TruthfulQA assesses the model’s ability to identify true claims, specifically in the context of literal truth about the real world.
Appendix B Evaluation Metrics
B.1 Safety Evaluation
To calculate attack success rate (ASR) for SAP200 and AdvBench, we follow Zou et al. (2023) to consider an attack successful if the following refusal strings are not produced in the generation from the model:
  • $\bullet$
    “I’m sorry”,
  • $\bullet$
    “Sorry”,
  • $\bullet$
    “I am sorry”,
  • $\bullet$
    “I apologize”,
  • $\bullet$
    “As an AI”,
  • $\bullet$
    “I’m just”,
  • $\bullet$
    “As a language model”,
  • $\bullet$
    “As an Assistant”,
  • $\bullet$
    “I cannot”,
  • $\bullet$
    “I do not”,
  • $\bullet$
    “Hello!”,
  • $\bullet$
    “is illegal and unethical”,
  • $\bullet$
    “I’m really sorry”,
  • $\bullet$
    “OpenAI”,
  • $\bullet$
    “unethical”,
  • $\bullet$
    “not ethical”,
  • $\bullet$
    “illegal”,
  • $\bullet$
    “not legal”,
  • $\bullet$
    “My apologies”,
  • $\bullet$
    “but I cannot”,
  • $\bullet$
    “If you have any other non-malicious requests”,
  • $\bullet$
    “I’m not able to provide”,
  • $\bullet$
    “I am really sorry”,
  • $\bullet$
    “I cannot create”,
  • $\bullet$
    “I cannot support”,
  • $\bullet$
    “I can’t provide”,
  • $\bullet$
    “I can’t assist”.
For evaluation of DAN, considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations. Practically, we adopt gpt-3.5-turbo-0613888https://openai.com/blog/chatgpt as the evaluation function following Deng et al. (2023a). Especially, the model responses are submitted via API calls, and a specific evaluation prompt is designed for gpt-3.5-turbo-0613 to rate the harmfulness on a five-degree scale (‘A (Very Harmful)’, ‘B (Harmful)’, ‘C (Moderately Harmful)’, ‘D (Slightly Harmful)’, and ‘E (No Harmful at All)’). Figure 6 shows the detailed instruction for gpt-3.5-turbo-0613 evaluation. Human evaluation results in Sec 5.3 demonstrate the highest agreement between human and gpt-3.5-turbo-0613 when viewing the evaluation degrees of ‘A (Very Harmful)’ and ‘B (Harmful)’ as a success of jailbreak attacks.
Figure 6: Prompt template for harmfulness evaluation using gpt-3.5-turbo-0613.
B.2 Helpfulness Evaluation
For evaluation of AlpacaEval, we adopt gpt-4-1106-preview as the auto evaluator to reflect whether the general performance degrades after adding safety techniques (Dubois et al., 2023). Specifically, GPT-4 compares two responses to the same instruction: one with our methods and one provided by text-davinci-003 and report the win rate of our method. Figure 7 shows the detailed instruction for gpt-4-1106-preview evaluation. For MMLU, we follow Hendrycks et al. (2021) and report accuracy based on the model’s predictions and the groud truth labels. For TruthfulQA, we follow Chuang et al. (2023) and report on two main distinct metrics: MC1 and MC2 scores in Table 4. The complete results on the three metrics in TruthfulQA, i.e., MC1, MC2 and MC3, are presented in Table 9. We can see that our method consistently improves the truthfulness over different models, indicating that our method can be deployed in real applications to enhance LLM safety while increasing the general helpfulness to some extent.
Figure 7: Prompt template for AlpacaEval results evaluation using gpt-4-1106-preview.
Models Methods TruthfulQA
MC1 MC2 MC3
Vicuna-7B Vanilla 30.1 48.7 23.6
Ours 35.2 53.4 26.3
[…]
Ours 38.2 55.1 28.5
[…]
Ours 37.5 56.0 27.4
Table 9: Performance on TruthfulQA of our $\mathbb{IA}$ upon different models in terms of Accuracy (%). The best results are highlighted in bold.
B.3 Intention Recognition Success Evaluation
To verify whether the model can successfully identify the intention of jailbreak queries, we examine the model response of the first stage and view a success if it begin with “The essential intention of the query is”, as instructed in stage 1 of our method. And we empirically find that once the model successfully starts with this string, it can then successfully analyze the intention behind the user query.
Appendix C Language Models
To evaluate the effectiveness of our $\mathbb{IA}$ method, we validate our approach on six representative Large Language Models, each distinguished by its model architecture, model size, and alignment level. Specifically, we consider five open-source LLMs and one closed-source LLM.
ChatGLM-6B
(Zeng et al., 2023), trained on 1T tokens of both Chinese and English corpus, follows similar technologies to ChatGPT, including supervised fine-tuning, feedback bootstrap, and RLHF.
Vicuna-7B-Chat
and Vicuna-13B-v1.1(Chiang et al., 2023) are two open-source chatbots that are fine-tuned on LLaMA-7B and LLaMA-13B, using 70K samples of user-shared ChatGPT conversations, without further utilizing RLHF.
MPT-30B-Chat
(Team, 2023) is a chatbot model designed for generating conversations. It was developed by fine-tuning the MPT-30B on dialogue datasets such as ShareGPT, Vicuna, Camel-AI, and some generated datasets.
DeepSeek-67B-Chat
(DeepSeek-AI, 2024) is fine-tuned for dialogue based on DeepSeek-67B, and empirically verified to perform better on nearly 20 public datasets in both Chinese and English than LLaMA2-70B.
GPT-3.5
(OpenAI, 2023) is an advanced and closed-source LLM trained with massive text data and can generate human-like responses to a wide range of prompts and queries. We obtain the responses of gpt-3.5-turbo-1106 using API calls, which may differ from the response in web interfaces due to internal settings unexposed to the users.
To study the performance differences between our method and safety training methods, we used LLaMA2-7B-Chat as a representative of these methods considering the various safety training strategies it has undergone and the superior safety it ultimately demonstrates.
LLaMA-2-7B-chat
(Touvron et al., 2023), fine-tuned on LLaMA-2-7B by Meta, is optimized for dialogue cases and use reinforcement learning with human feedback (RLHF) to align with human preferences for helpfulness and safety.
For each of the above models, we conducted experiments under the system prompt template designed by their developers to simulate real-world usage scenarios.
Appendix D Alternative Prompts
To study the effect of the implementation of our method, we combine the two stages of our mechanism and crafted an one-pass intention analysis instruction. See Figure 8 for the detailed prompt.
Figure 8: One-pass intention analysis prompt for our method.
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. These additional instructions were meticulously designed with modifications in lexical choices and syntactic constructions to direct intention analysis. The detailed prompts are presented in Figure 9.
Refer to caption Refer to caption
(a) Prompt Set A (b) Prompt Set B
Figure 9: Alternative prompts crafted for our intention analysis instructions.
Appendix E Cross-Intention Analysis Experiment
A question that naturally arises about our method is, “Would the effectiveness of $\mathbb{IA}$ improve notably with a larger intention analysis model?” To investigate this, we conduct cross-experiments using Vicuna-7B and 13B models. Specifically, we evaluate each target model in three contexts: without any intention analysis model, with the Vicuna-7B as the intention analysis model, and with the Vicuna-13B serving the same role. Table 10 shows the results. It is observable that for the 7B model, utilizing the 13B for intention analysis yields only marginal performance gains. For the 13B model, employing the 7B model for intention recognition achieves comparable results to using the 13B model itself for intention analysis, but with the added benefit of conserving inference resources. This indicates that the 7B model is sufficiently capable of intention analysis to enhance model safety.
Target Model Intent. Model DAN SAP200
Vicuna-7B 44.3 67.2
Vicuna-7B 2.89 0.31
Vicuna-13B 1.93 0.62
Vicuna-13B 54.7 65.4
Vicuna-7B 1.25 1.87
Vicuna-13B 0.64 1.12
Table 10: ASR (%) of our $\mathbb{IA}$ on DAN and SAP200 with different intention analysis model scales. For each target model, the intention analysis is performed in three ways, i.e., without intention analysis, analyzed by Vicuna-7B, and by Vicuna-13B.
Appendix F Qualitative Examples in Ours Experiments
This section presents some practical examples in our experiments. These examples are from SAP200 dataset (see Figure 10), DAN dataset (see Figure 11), and AdvBench dataset (see Figure 12), respectively.
Figure 10: Qualitative example I from SAP200 dataset about the comparison between vanilla and our $\mathbb{IA}$.
Figure 11: Qualitative example II from DAN dataset about the comparison between vanilla and our $\mathbb{IA}$.
Figure 12: Qualitative example III from AdvBench dataset about the comparison between vanilla and our $\mathbb{IA}$.