### v4

# Intention Analysis Makes LLMs A Good Jailbreak Defender

Abstract
Aligning large language models (LLMs) with human values, particularly when facing complex and stealthy jailbreak attacks, presents a formidable challenge. Unfortunately, existing methods often overlook this intrinsic nature of jailbreaks, which limits their effectiveness in such complex scenarios. In this study, we present a simple yet highly effective defense strategy, i.e., Intention Analysis ($\mathbb{IA}$). $\mathbb{IA}$ works by triggering LLMs’ inherent self-correct and improve ability through a two-stage process: 1) analyzing the essential intention of the user input, and 2) providing final policy-aligned responses based on the first round conversation. Notably, $\mathbb{IA}$ is an inference-only method, thus could enhance LLM safety without compromising their helpfulness [^1] Improving the safety of LLM with training-required methods Ouyang et al. (2022); Touvron et al. (2023) always necessitates great effort to strike a delicate balance between safety and helpfulness.. Extensive experiments on varying jailbreak benchmarks across a wide range of LLMs show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -48.2% attack success rate). Encouragingly, with our $\mathbb{IA}$, Vicuna-7B even outperforms GPT-3.5 regarding attack success rate. We empirically demonstrate that, to some extent, $\mathbb{IA}$ is robust to errors in generated intentions. Further analyses reveal the underlying principle of $\mathbb{IA}$: suppressing LLM’s tendency to follow jailbreak prompts, thereby enhancing safety.
Warning: Some of the examples may be harmful!
Intention Analysis Makes LLMs A Good Jailbreak Defender
Yuqi Zhang1, Liang Ding2, Lefei Zhang1thanks:  Corresponding Author., Dacheng Tao3 1School of Computer Science, Wuhan University   2The University of Sydney 3College of Computing and Data Science at Nanyang Technological University, Singapore 639798 [Uncaptioned image] {zhangyuqi,zhanglefei}@whu.edu.cn,  {liangding.liam,dacheng.tao}@gmail.com [Uncaptioned image] https://github.com/alphadl/SafeLLM_with_IntentionAnalysis
1 Introduction
Recently, Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2023; Google, 2023), such as ChatGPT, not only show remarkable capabilities in various tasks Qin et al. (2023); Zhong et al. (2023); Peng et al. (2023); Ren et al. (2024), but also lead to the risk of potential misuse (e.g., producing harmful responses or illegal suggestions) Weidinger et al. (2021). Efforts like Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) have been made to alleviate these risks and enhance LLMs’ alignment with human values, making LLMs able to refuse direct harmful questions like how to rob a bank? However, LLMs remain vulnerable to some adversarial inputs, particularly in the context of so-called “jailbreak” attacks. These jailbreaks are specially designed to circumvent safety policy and manipulate LLMs for their restricted outputs (Yuan et al., 2024; Zou et al., 2023), which poses formidable risks in real applications.
Figure 1: Performance of our method on different LLMs. Our $\mathbb{IA}$ 1) reduces Attack Success Rate ($\downarrow$) against both crafted jailbreak prompts (DAN and DeepInception) and automatic attack (GCG), 2) achieves remarkable safety improvements for both SFT (Vicuna-7B & MPT-30B-Chat) and RLHF (GPT-3.5) LLMs.
To defend LLMs against jailbreak attacks, existing popular methods either focus on emphasizing safety during inference (Xie et al., 2023; Wei et al., 2023b), or modifying the user inputs (Robey et al., 2023) or evaluating inputs/outputs’ safety (Li et al., 2024), often neglecting the intrinsic characteristics of jailbreak attacks. This oversight limits their effectiveness in more complex jailbreak scenarios (see experimental results in Section 4.2). Through analysis, we find that these jailbreaks typically work by concealing harmful questions within seemingly inoffensive and complex scenarios, such as role-playing or virtual scene construction (Liu et al., 2023b). Such disguise leads LLMs to focus on the jailbreak prompt excessively, impairing their awareness of the harmful question itself (See Figure 5 for evidence). We assume such insufficient awareness of the harmful content concealed in complex jailbreak queries is the fundamental reason for LLM’s vulnerability to these attacks. Drawing insights from classic dialogue system design (Allen and Perrault, 1980), an effective solution is to tailor an intent recognition mechanism specifically for jailbreak scenarios to enhance LLM’s understanding of user queries regarding safety and improve its ability to recognize concealed harmful questions.
In this paper, we leverage the intrinsic intent recognition capabilities of LLMs, proposing an Intention Analysis ($\mathbb{IA}$) strategy. Specifically, $\mathbb{IA}$ enables LLMs to analyze the essential intention of the user query to better understand it and recognize the underlying unsafe content within before finally responding, as shown in Figure 2. Such intention analysis mechanism can significantly improve LLM safety against varying jailbreak attacks, see Figure 1 for a demonstration. We dive deeper from the perspective of attention scores and find that the underlying principle of $\mathbb{IA}$ is to suppress LLM’s tendency to follow jailbreak prompts. Notably, our $\mathbb{IA}$ is an inference-only method that can significantly enhance LLM safety without the need for additional safety training (Ouyang et al., 2022; Touvron et al., 2023). In this way, $\mathbb{IA}$ skillfully circumvents the safety-helpfulness trade-off and enables comparable safety improvement as well as better helpfulness maintenance.
To summarize, our contributions are as follows:
  • $\bullet$
    We introduce $\mathbb{IA}$, a new method that significantly enhances LLM safety in the context of sophisticated jailbreak attacks through an intention analysis mechanism.
  • $\bullet$
    $\mathbb{IA}$ is a plug-and-play inference-only method, thereby 1) cleverly circumventing the safety-helpfulness trade-off that is challenging in safety training, and 2) can be flexibly and effectively deployed upon any LLMs.
  • $\bullet$
    Empirically, our robust $\mathbb{IA}$ significantly and consistently reduces the harmfulness of LLM outputs, while maintaining the helpfulness, achieving new state-of-the-art performance on several benchmarks, e.g., DeepInception.
Figure 2: Illustrated Comparison of (a) vanilla and (b) the proposed $\mathbb{IA}$. Our $\mathbb{IA}$ consists of two stages: (1) Essential Intention Analysis: instructing the language model to analyse the intention of the user query with an emphasis on safety, ethics, and legality; (2) Policy-Aligned Response: eliciting the final response aligned with safety policy, building upon the analyzed intention from the first stage.
2 Related Work
Alignment-Breaking Adversarial Attack
Despite significant efforts to align LLMs with human preference (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Korbak et al., 2023; Miao et al., 2024), adversarial attackers can still elicit harmful responses from LLMs by “jailbreak” attacks (Shen et al., 2023; Liu et al., 2023b). Current jailbreak attack methods are primarily classified into two categories: in-the-wild jailbreak prompts and optimization-based automatic attacks (Chao et al., 2023; Yu et al., 2023). In-the-wild jailbreak prompts are typically hand-crafted through human ingenuity and is semantically understandable in general Shen et al. (2023). For optimization-based automatic attacks, a representative work is to automatically fetch a transferable attack suffix through the Greedy Coordinate Gradient (GCG) algorithm which maximizes the probability of the language model generating an affirmative and unsafe response Zou et al. (2023). In this work, various attacks mentioned above are considered in experiments to comprehensively test the defensive performance of our method.
Safety-Enhancing Defensive Methods
Recently, numerous methods have been developed to reduce LLMs’ harmful generations in inference stage. A branch of them mainly concentrates on controlling the content that LLMs can see by pre-processing user inputs, such as perplexity filtering (Alon and Kamfonas, 2023; Jain et al., 2023), paraphrasing (Jain et al., 2023) and re-tokenization (Cao et al., 2023; Jain et al., 2023). Another branch focuses on exploiting LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks, such as letting LLMs self-evaluate their outputs (Helbling et al., 2023; Li et al., 2024; Wang et al., 2024) or reminding of safety in system mode with conventional decoding Xie et al. (2023) or contrastive decoding Zhong et al. (2024).
While existing methods effectively prevent unsafe responses, their efficacy drops significantly against sophisticated jailbreak attacks that conceal harmful questions within complex and seemingly inoffensive scenarios. In contrast, our method enhances LLM safety by leveraging the intrinsic intent recognition capabilities of LLMs to detect these concealed threats (see Table 1 for details).
3 Methodology
3.1 Preliminary
We focus on enhancing LLM safety during the inference stage. In practice, developers usually implement pre-defined system prompts for LLMs to facilitate safe, responsible, and effective interactions with users (Chiang et al., 2023). Under this premise, the system prompt $P_{sys}$ and the user prompt $P_{usr}$ are concatenated to form the final input $\{x_{1:n}^{s},x_{1:m}^{u}\}$ of the LLM, where $P_{sys}=\{x_{1}^{s},x_{2}^{s},\dots,x_{n}^{s}\}$, $P_{usr}=\{x_{1}^{u},x_{2}^{u},\dots,x_{m}^{u}\}$, $x_{i}^{s}$ and $x_{j}^{u}$ are the $i$-th and $j$-th token of $P_{sys}$ and $P_{usr}$, respectively. Conditioned on the input $\{x_{1:n}^{s},x_{1:m}^{u}\}$, the autoregressive inference process of response $R=y_{1:L}$ is formulated as following:
$\vspace{-0.2cm}q(y_{1:L}|x_{1:n}^{s},x_{1:m}^{u})=\prod\limits_{i=1}^{L}q(y_{i}|y_{1:i-1},x_{1:n}^{s},x_{1:m}^{u}).\vspace{-0.2cm}$
For simplicity, we use $R\sim q(R|P_{sys},P_{usr)}$ to denote sampling a response $R$ from $q(\cdot)$ given the prompt $P_{sys}$ and $P_{usr}$. In this way, the response $R$ can be obtained as: $R=\textrm{LLM}\left(P_{sys},P_{usr}\right).$
In this work, we aim to leverage LLMs’ intrinsic abilities of intention analysis, to enhance their safety against varying jailbreak attacks during the inference stage, while simultaneously maintaining the general helpfulness.
3.2 $\mathbb{IA}$: Intention Analysis
To achieve the above goal, we introduce $\mathbb{IA}$, a zero-shot intention analysis mechanism, to guide LLMs to explicitly identify and understand the underlying intention of a user query before facilitate a final response. Specifically, we devise a two-stage intention analysis instruction to accomplish the whole process [^2] Full prompts can be found in Figure 9., as illustrated in Figure 2(b): (1) essential intention analysis and (2) policy-aligned response.
Stage 1: Essential Intention Analysis
This stage focuses on guiding the LLMs to discern the core intention behind the user query, with a specific orientation towards safety, ethics, and legality. The critical question arises: How can we ensure that LLMs accurately identify the query’s intention? Actually, recent studies Bender and Koller (2020); Zhu et al. (2024); Gómez-Pérez et al. (2023) have shown that LLMs are notably proficient at language understanding tasks, and intention analysis is a straightforward task, indicating the competence of LLMs in performing this stage. The only concern is generative models’ potential hallucination when performing the discriminative tasks Ji et al. (2023); Yan et al. (2021); Ye et al. (2023); Lu et al. (2024), therefore, we carefully define the format for the models’ response, that is, beginning with “The essential intention of the query is”, which has been validated in our analysis.
In practice, we construct the instruction for the LLMs to effectively perform intention analysis, denoted as $I_{rec}$. When presented with a user query $P_{usr}$ [^3] In this context, the user query $P_{usr}$ mostly represents the entirety of a jailbreak query., we concatenated $I_{rec}$ and $P_{usr}$ to form a whole “User” level input $I_{rec}\oplus P_{usr}$ for the LLMs. Subsequently, the designated target LLMs engage in an auto-regressive inference process, guided by its system prompt $P_{sys}$, to produce the stage-specific response:
$R_{st1}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr}\right),\vspace{-0.3cm}$
which is expected to contain the essential intention.
Stage 2: Policy-Aligned Response
Having successfully recognized the essential intention, the second stage aims to elicit the final response from the LLMs. We first direct the LLMs to bear the identified intention in mind and then provide a final response to the user query. Meanwhile, we explicitly instruct the LLMs to strictly adhere to safety policy and ethical standards [^4] The details of safety policy and ethical standards are not explicitly provided because we have found they significantly increase inference costs with minimal benefit. We believe that LLMs, through training, develop an inherent understanding of safety, allowing implicit prompts to effectively activate this internal knowledge. and ensure the exclusion of any unsafe content in their responses. To this end, the second stage further strengthens the role of the intention analysis and reinforces the inherent alignment of LLMs with the safety policy.
Specifically, we concatenate the dialogue from the first stage with the instruction for the current stage, denoted as $I_{ct}$, forming the complete input for LLMs. Then a similar autoregressive inference process is conducted, leading to the generation of the final response $R_{st2}$ to the user query $P_{usr}$:
$R_{st2}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr},R_{st1},I_{ct}\right).\vspace{-0.2cm}$
To assess the safety of the response, we follow Shen et al. (2023) to employ a binary auto-annotation function $\textrm{AS}(\cdot)$ [^5] Will be discussed in detail in Section 4.1. to determine the harmfulness of $R_{st2}$. If the outcome yields $\textrm{AS}(R_{st2})=False,$ then the response is deemed safe, indicating a successful defense against the jailbreak attack.
Models Defense Methods Attack Methods Average Time Cost
DAN SAP200 DeepInception GCG AutoDAN
ChatGLM-6B Vanilla 29.0 45.8 100 88.0 99.5 72.5 14.3
Input Check 16.3 9.52 46.2 9.00 89.0 34.0 12.6
ICD 19.1 2.81 17.1 17.0 2.00 11.6 15.2
Self-Reminder 22.5 3.13 17.9 0.00 66.0 21.9 17.1
SmoothLLM 7.19 20.6 84.5 1.00 84.0 39.5 113.4
$\mathbb{IA}$ (Ours) 5.48 6.12 0.00 1.00 2.00 2.92 19.2
[…]
Input Check 7.50 0.00 0.00 0.00 43.0 10.1 10.7
ICD 0.98 0.00 0.00 0.00 0.00 0.20 15.5
Self-Reminder 0.77 0.00 4.38 0.00 0.00 1.03 14.8
SmoothLLM 0.31 2.81 86.5 0.00 71.5 32.2 118.5
$\mathbb{IA}$ (Ours) 0.13 0.00 0.00 0.00 0.00 0.03 19.5
[…]
Input Check 3.43 0.00 7.57 0.00 7.00 3.60 4.98
ICD 0.63 0.00 0.00 0.00 0.00 0.13 5.12
Self-Reminder 0.63 0.00 0.00 0.00 0.00 0.13 6.64
SmoothLLM 0.31 0.63 32.7 0.00 46.0 15.9 79.2
$\mathbb{IA}$ (Ours) 0.31 0.00 0.00 0.00 0.00 0.06 10.6
[…]
Input Check 19.0 58.1 53.0 13.0 100 48.6 8.64
ICD 40.4 32.8 0.00 1.00 88.0 32.4 10.3
Self-Reminder 41.3 33.8 55.4 11.0 98.5 48.0 15.0
SmoothLLM 13.5 54.4 96.4 8.00 98.5 54.2 102.7
$\mathbb{IA}$ (Ours) 3.42 0.31 0.00 0.00 10.5 2.85 17.3
[…]
Input Check 7.19 7.50 98.8 3.00 97.5 42.8 10.7
ICD 53.9 32.8 86.9 0.00 91.5 53.0 13.1
Self-Reminder 52.5 36.9 75.7 1.00 83.0 49.8 16.4
SmoothLLM 17.3 37.0 94.0 5.00 98.0 50.3 136.1
$\mathbb{IA}$ (Ours) 0.94 1.12 0.00 0.00 3.50 1.11 22.1
[…]
Input Check 14.1 9.38 41.8 6.00 14.3 132.2
ICD 49.4 29.9 100.0 3.00 45.6 218.7
Self-Reminder 46.9 39.4 100 19.0 51.3 210.0
SmoothLLM 60.6 64.4 22.0 22.0 42.3 534.8
$\mathbb{IA}$ (Ours) 5.38 19.2 4.78 4.00 8.34 223.0
[…]
Input Check 30.3 3.20 5.80 1.00 8.06 154.2
ICD 45.6 14.4 47.8 9.00 29.2 162.8
Self-Reminder 9.58 7.81 3.19 1.00 5.40 177.4
SmoothLLM 26.9 11.9 51.0 0.00 22.4 486.6
$\mathbb{IA}$ (Ours) 3.78 1.56 7.57 2.00 3.73 198.0
[…]
Input Check 2.50 0.00 0.00 4.00 1.63 2.47
ICD 0.94 0.31 0.00 0.00 0.31 5.12
Self-Reminder 2.81 0.31 0.00 0.00 0.78 7.21
SmoothLLM 0.64 0.00 0.00 0.00 0.16 15.2
$\mathbb{IA}$ (Ours) 0.64 0.00 0.00 0.00 0.16 8.27
Table 1: Comparison of our $\mathbb{IA}$ and four baselines under five jailbreak methods in terms of ASR (%) and time cost (s/sample). The best and second best average results are highlighted in bold and underline . Among them, DAN, SAP200, and DeepInception are complex and stealthy in-the-wild jailbreaks, while GCG and AutoDAN are optimization-based automatic jailbreaks. “—” means lacking official AutoDAN implementation for distributed larger models (MPT-30B and DeekSeek-67B) or white-box LLM weights required (GPT-3.5).
4 Experiment
4.1 Setup
Datasets
For safety datasets, we experiment on three main categories of jailbreak attacks, including three representative complex and stealthy in-the-wild jailbreak datasets (i.e. DAN (Shen et al., 2023), SAP200 (Deng et al., 2023a), and DeepInception (Li et al., 2023)), two popular optimization-based automatic jailbreak methods (i.e. GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023a)), and two advanced attacks for GPT-3.5 (i.e. multilingual attack called MultiJail (Deng et al., 2023b) and encryption-based attack named CipherChat (Yuan et al., 2024)).
Besides, to evaluate $\mathbb{IA}$’s effect on helpfulness for general benign queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022).
Evaluation Metrics
For safety assessment, we annotate the harmfulness of responses and report attack success rate (ASR, Shen et al., 2023), where lower scores indicate stronger safety. Specifically, for DAN dataset, considering the complexity of responses, we adopt gpt-3.5-turbo-0613666https://openai.com/blog/chatgpt as the auto-annotation function following Deng et al. (2023a) and carry our human evaluation in Appendix C.1 to ensure the credibility. For other safety datasets, we annotate harmfulness following Zou et al. (2023) by matching refusal strings (e.g., “I’m sorry”; see Appendix C.2 for detailed settings).
For helpfulness assessment, we report win rate (Dubois et al., 2023) for AlpacaEval and accuracy (Hendrycks et al., 2021) for MMLU. For TruthfulQA, we follow Chuang et al. (2023) and report on two distinct metrics: MC1 and MC2 scores, where higher scores indicate stronger factuality (see Appendix C.3 for more details).
Models
To evaluate $\mathbb{IA}$’s effectiveness, we experiment on representative LLMs with varying scales and alignment levels, including not only SFT models, i.e. Vicuna-7B/13B-v1.1 (Chiang et al., 2023) and MPT-30B-Chat (Team, 2023), but also RLHF models, i.e. ChatGLM-6B (Zeng et al., 2023), Llama2-7B-Chat (Touvron et al., 2023), Llama3-8B-Instruct [^7] https://ai.meta.com/blog/meta-Llama-3/, and DeepSeek-67B-Chat (DeepSeek-AI, 2024). Beyond open-source LLMs, our experimentation extends to an advanced closed-source LLM, GPT-3.5 (gpt-3.5-turbo-1106) (OpenAI, 2023), renowned for its superior capabilities, especially safety alignment.
Comparison Baselines
We compare our $\mathbb{IA}$ with vanilla LLMs (no defense) and seven popular defense methods, i.e., Input Check [^8] We create an Input Check baseline by using the prompt in Helbling et al. (2023) and operate in the input space to let LLMs judge whether a query is harmful or not., ICD (Wei et al., 2023b), (System-Mode) Self-Reminder (Xie et al., 2023), SmoothLLM (Robey et al., 2023), BPE-dropout (Jain et al., 2023), Self Defense (Helbling et al., 2023), and Moral Self-Correction (Ganguli et al., 2022). The first four representative defense methods are reported in Table 1 and others in Table 7 in Appendix due to page limitation. Besides, a training method is also included in Appendix E.3 and results show $\mathbb{IA}$ achieves both safety and helpfulness goals without additional resource-consuming safety training. For a fair comparison, we closely follow the best default parameters in their papers.
Implementation
The detailed $\mathbb{IA}$ prompts for experiments are provided in Figure $9$To assess the resilience of our method against specific expressions, we construct two alternative $\mathbb{IA}$ prompts in Appendix D.2 and experiment results demonstrate $\mathbb{IA}$’s effectiveness is irrespective of specific expressions.. For the DAN dataset, we compile an evaluation dataset of 1560 samples by extracting 195 questions from each jailbreak community within the forbidden question set (Shen et al., 2023). For GCG, we follow Zou et al. (2023) and conduct transfer attacks on Vicuna-7B and 13B. The adversarial suffix achieving the lowest loss after 500 steps of optimization are adopted to further attack target models on 100 individual harmful behaviors (Wei et al., 2023b). For open-source models, we download them from HuggingFace [^10] https://huggingface.co/models. For closed-source models, we obtain the responses of GPT-3.5 via API calls. Throughout our experiments, we set a temperature of zero for deterministic outcomes Peng et al. (2023) and a generation length of 1024 tokens, employing default system prompt templates for each LLM regarding their official reports. All experiments are carried out on a solitary node outfitted with 8 A100-SXM80GB GPUs.
Figure 3: The confusion matrix illustrating the relationship between the success of intention analysis and the harmlessness of LLM’s final response on SAP200 and DAN datasets. “IR Succ.” and “IR Fail.” represent success or failure of intention analysis, respectively.
Models Methods AlpacaEval MMLU TruthfulQA
Win Rate Acc. MC1 MC2
[…]
Table 2: General performance on helpful dataset upon different models in terms of Win Rate (%) for AlpacaEval, Accuracy (%) for MMLU and MC1, MC2 (%) for TruthfulQA. “—” means lacking official implementation for distributed larger models or white-box LLM weights required.
4.2 Main Results
Performance of safety on various jailbreak attacks
In Table 1, we represent the ASR of several defense baselines on different LLMs under various jailbreak attacks as well as inference time comparison [^11] Due to memory constraints, the Deepspeed Zero-3 algorithm was employed for larger models, MPT-30B and DeepSeek-67B, resulting in relatively longer inference times.. We can observe that: 1) $\mathbb{IA}$ effectively reduces ASRs across a diverse range of LLMs along with an acceptable time cost. For LLMs with high vanilla ASRs, such as ChatGLM-6B, Vicuna-7B, Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat, we significantly lower the average ASRs from 72.7% to 3.79%. Similarly, for LLMs presenting lower vanilla ASRs, such as Llama2-7B-Chat, Llama3-8B-Instruct, and GPT-3.5, $\mathbb{IA}$ further reduces their average ASRs from 13.8% to mere 0.1%. 2) $\mathbb{IA}$ maintain its effectiveness even in scenarios where other defense methods struggle. For example, AutoDAN leverages LLMs to automatically attack based on optimization and thus is hard to defend. While the baselines have ASRs of at least 83% on Vicuna-7B and 13B under AutoDAN, $\mathbb{IA}$ can still provide significant defense with a low ASR of under 11%. Notably, $\mathbb{IA}$ can also integrate with another defensive method to enhance performance but with additional computation overhead (see Appendix E.4 for details). Moreover, we also extend to more advanced jailbreak attacks including multilingual and encryption-based attacks, and demonstrate our consistent effectiveness on ChatGPT (see Appendix E.1). Further analysis regarding our good performance will be discussed in Section 5.
Performance of general helpfulness for benign queries
An effective defense method is expected to maintain general abilities as well. To explore the impact of our method on the general performance of LLMs, we conduct experiments on several acknowledged helpfulness datasets and report the results in Table 2. As observed, for harmless user prompts, our $\mathbb{IA}$ does not significantly compromise the general helpfulness on AlpacaEval, MMLU, and TruthfulQA benchmarks compared with vanilla LLMs. These results indicate that $\mathbb{IA}$ can be deployed in real applications to enhance LLM safety while preserving general helpfulness. More comparison results with other defensive methods can be found in Table 11 in Appendix. To further study $\mathbb{IA}$’s impact on LLM’s helpfulness, we also conduct both manual and automatic checks about safe refusal’s helpfulness for harmful queries and find that $\mathbb{IA}$ enables LLMs to effectively give safe refusals with satisfactory helpfulness for harmful queries, instead of simple rejection (see Appendix G for detailed analysis).
5 Discussion of $\mathbb{IA}$ Mechanism
5.1 Can LLMs successfully generate the intentions behind jailbreak queries?
Intention analysis is a straightforward language understanding task for LLMs to proficiently perform (Bender and Koller, 2020; Zhu et al., 2024; Gómez-Pérez et al., 2023). The results of intention analysis are binary—either LLMs can successfully detect the intention, such as identifying plans to “rob a bank” as shown in Figure 2, or they fail and miss it. In Figure 3, we count the samples and examine the correlation between successful intention analysis (see Appendix C.4 for evaluation details) and producing harmless responses on SAP200 and DAN datasets [^12] SAP200 and DAN datasets are chosen for intention analysis evaluation due to their most complex and stealthy intentions among jailbreak datasets tested..
We observe that: 1) Most LLMs are highly effective in recognizing intentions behind complex and stealthy jailbreak queries, achieving a nearly 100% success rate in Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat. Particularly, the intention recognition rate of Llama2-7B-Chat is relatively lower due to its excessively strong inherent safety leading to direct refusals to harmful user queries [^13] We do not present Llama3-8B-Instruct for the same reason that its strong inherent safety leads to almost all direct refusals in the intention analysis stage.(see Figure 17 for detailed cases). 2) In adversarial scenarios, it is easier for most LLMs to generate intentions than directly generate safe responses. Setting the SAP dataset as an example, most LLMs can successfully identify more than 90% of the adversarial intents. While in Table 1, they can only generate averagely $\sim$30safe responses.
Figure 4: Performance of $\mathbb{IA}$ with varying correct intention ratio on DAN dataset. From left to right: the correct intentions are replaced with masked and random intention, respectively.
5.2 What if LLMs generate incorrect intentions?
To understand the effect of intention analysis errors, we examine two extreme cases: 1) recognized intentions are masked with an invalid field (e.g., “[secret]”); 2) recognized intentions are replaced with randomly sampled tokens from LLM’s vocabulary, simulating a severely wrong case. Figure 4 shows $\mathbb{IA}$’s performance across different correct intention ratios on DAN dataset. Overall, $\mathbb{IA}$’s performance declines with increasing intention errors but consistently maintains a much lower ASR (below 10%) compared to the vanilla baseline, indicating $\mathbb{IA}$’s some robustness to wrong intentions.
Notably, $\mathbb{IA}$ remains effective even at a 0% correct intention ratio. This can be attributed to the role of the intention analysis sequence format, allowing replacing true intentions with invalid ones to be marginally detrimental, as widely recognized by the In Context Learning (ICL) community (Min et al., 2022). Further explanation can be found at Appendix F. However, exploring the underlying principles of how sequence formats affect outcomes is beyond the scope of this work.
5.3 What is the underlying principle of $\mathbb{IA}$?
This section explores how $\mathbb{IA}$ works by analyzing the model’s attention distribution across different prompt components during response generation [^14] Inspired by Wang et al. (2023), the attention score is calculated by averaging the maximum attention scores for each prompt component across the samples in the DAN dataset. (see Figure 5) [^15] As depicted in Figure 2, vanilla prompt consists of jailbreak prompt and harmful question. IA-Stage 1 prompt adds an intention analysis instruction before the vanilla prompt. IA-Stage 2 then combines the IA-Stage 1 prompt, the recognized intention from Stage 1, and the final response instruction.. As shown, the model under vanilla prompt pays significant attention to the jailbreak prompt, leading to potentially harmful responses. In contrast, $\mathbb{IA}$ at both stages significantly reduces LLM’s attention to the jailbreak prompt while increasing attention to user intent, making LLM less likely to follow jailbreak prompts and leading to safer responses.
To further illustrate $\mathbb{IA}$’s effect, Figure 6 presents a layer-wise comparison of attention on the jailbreak prompt between the vanilla and $\mathbb{IA}$ prompts. The results show that $\mathbb{IA}$ consistently reduces the model’s attention on the jailbreak prompt across all layers, further indicating $\mathbb{IA}$’s effectiveness in suppressing LLM’s tendency to follow jailbreak prompts.
Figure 5: Comparison of Vicuna-13B’s attention scores on different prompt components of different methods. The average attention score is computed on DAN dataset. $\mathbb{IA}$ largely decreases model’s attention to jailbreak prompt (red bar) in both two stages.
6 Further Discussion
Two factors influence $\mathbb{IA}$ performance. (1) Intention analysis ability: As shown by the solid lines in Figure 4, $\mathbb{IA}$ performance improves with higher correct intention ratios, suggesting that better intention analysis ability can further enhance effectiveness [^16] We also conduct cross-intention analysis experiment on Vicuna-7B and Vicuna-13B in Appendix H to explore the effect of different intention analysis LLMs.. (2) Inherent LLM safety: Figure 3 shows that even among LLMs with nearly 100% intention recognition rates, the final harmful response rates vary notably—from 0.3% for Vicuna-7B to 19.3% for MPT-30B-Chat—highlighting the impact of inherent LLM safety on $\mathbb{IA}$ results (see Figure 18 for a related case study). These suggest two improvement directions: enhancing LLMs’ intention analysis ability and their inherent safety.
Our efficient one-pass variant of $\mathbb{IA}$ provides a more cost-effective choice.
As aforementioned, to maximize the performance, our $\mathbb{IA}$ follows a two-stage process. A natural question arises of whether our mechanism can be merged into one step, to save the decoding overhead. To verify this, we design a cheaper one-pass $\mathbb{IA}$ variant (see Figure 10 for detailed prompts). From results in Table 3, we see that: 1) For more powerful models, such as Vicuna-7B and 13B, one-pass $\mathbb{IA}$ achieves comparable performance to two-stage $\mathbb{IA}$ in a more cost-effective manner. 2) For less powerful models, i.e., ChatGLM-6B, one-pass $\mathbb{IA}$’s effectiveness diminishes to some extent. In such cases, two-stage $\mathbb{IA}$ is necessary to sustain satisfactory performance.
Figure 6: Comparison of Vicuna-13B’s attention scores on jailbreak prompt between Vanilla and $\mathbb{IA}$ methods across different model layers. The average attention score is computed on DAN dataset. High scores means greater influence of jailbreak prompt on the generated response.
Vicuna-7B Vicuna-13B ChatGLM-6B Time Cost
Vanilla 73.4 65.4 45.8 13.2
+ One-Pass $\mathbb{IA}$ 5.50 1.13 39.0 12.6
+ Two-Stage $\mathbb{IA}$ 0.31 1.12 6.12 19.5
Table 3: Comparison of our $\mathbb{IA}$ with different implementations (one-pass and two-stage) on SAP200 in terms of ASR (%) and average Time Cost (s/sample).
7 Conclusion
In this work, a simple yet highly effective defense strategy $\mathbb{IA}$ is proposed to handle the widespread complex and stealthy jailbreak attacks. $\mathbb{IA}$ leverages LLM’s intrinsic capacities to analyze the essential intention of user queries before finally responding through two stages. Extensive experiments on representative jailbreak benchmarks across diverse LLMs show that $\mathbb{IA}$ could consistently and significantly enhance LLM safety while maintaining general helpfulness. $\mathbb{IA}$ works by suppressing LLM’s tendency to follow jailbreak prompts, thus leading to safer responses. Further analysis indicates that enhancing LLMs’ intention analysis capability and their inherent safety are two directions for future improvements.
Limitations
Our method remains to be validated on more advanced models. However, since our core intention analysis mechanism relies on LLM’s fundamental capabilities of—specifically, instruction-following and text comprehension—making it easy to perform, we believe this approach has the potential to generalize effectively across diverse models as a safety mechanism. Additionally, despite the effectiveness of our method in defending sophisticated jailbreak prompts, these prompts do not encompass the entire potential jailbreak attacks encountered in real-world scenarios. Consequently, the practical applicability of our approach remains to be validated through further testing. Our research underlines the importance of intention analysis in improving LLM safety, suggesting future work focusing on integrating this into training to reduce inference costs. Additionally, in the face of the rapid advancements in the adversarial attacks community, there is a pressing need for developing more effective and robust defense strategies for LLMs. While our method specifically targets jailbreak scenarios, broader alignment tasks still benefit from alignment training, such as RLHF.
Ethics Statement
We take ethical considerations very seriously. This paper focuses on improving the safety (especially the jailbreak attacks) of large language models, through carefully designed intention analysis prompting mechanism. Our research could significantly reduce the unsafe responses of LLMs. All experiments are conducted on open datasets and the findings and conclusions of this paper are reported accurately and objectively. Thus, we believe that this research will not pose ethical issues.
Acknowledgments
We express our gratitude to Zuchao Li and Yuchun Miao for their assistance with proofreading and insightful feedback on the writing of this paper. We thank the anonymous reviewers and the area chair for their insightful comments and suggestions. This research is supported by the National Research Foundation, Singapore, and the CyberSG R&D Programme Office (“CRPO”), under the National Cybersecurity R&D Programme (“NCRP”), RIE2025 NCRP Funding Initiative (Award CRPO-GC1-NTU-002).
Appendix A Experimental Datasets
A.1 Safety Datasets
Hand-Crafted Jailbreak Prompts
To assess the effectiveness of our method on in-the-wild jailbreak prompts, we employ two jailbreak prompt datasets. The first is forbidden question set developed by Shen et al. (2023), which is currently the largest in-the-wild jailbreak prompt dataset. To improve computing efficiency, we extract five questions from each forbidden scenario, forming a jailbreak dataset comprising 8 jailbreak communities $\times$ 3 jailbreak prompts $\times$ 13 forbidden scenarios $\times$ 5 questions, totaling 1560 samples. The term “DAN” is used to denote this dataset. For evaluation, we leverage attack success rate (ASR) to consider the success of a jailbreak attack. Considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations, as will be discussed in Section C.2.
The second SAP200 is an jailbreak prompt dataset, constructed semi-automatically by Deng et al. (2023a) using code injection and payload splitting mechanisms. It encompasses 8 distinct sensitive topics, with 200 samples each, totaling 1600 samples.
Due to computational resource and financial limitations, we randomly select 40 samples for each sub-dataset, totaling $32040samples\times 8sub-datasets=$ 320 samples from DAN and SAP200 datasets,respectively, to conduct comparative experiments in Table 7 and correct intention ratio comparison experiments in Figure 4.
Gradient-Based Adversarial Attacks
To comprehensively verify the effectiveness of our method in defending against jailbreak attacks, we conduct experiments on a popular token-level jailbreak dataset, i.e., AdvBench (Zou et al., 2023) and use the Greedy Coordinate Gradient (GCG) attack algorithm to generate the adversarial suffix. Specifically, we utilize Vicuna-7B and 13B to optimize a universal attack suffix by combining the gradients of the two models. Subsequently, we use the held-out 100 harmful behaviors from AdvBench and apply this optimized suffix to attack other models. We followed the same default parameter setting for GCG, with a learning rate of 0.01, batch size of 512, top-k of 256, and temperature of 1. The suffix achieving the lowest loss after 500 steps was selected for the experiment.
A.2 Helpfulness Datasets
To evaluate the effect of our $\mathbb{IA}$ on helpfulness for general in-distribution queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022). AlpacaEval, containing 805 general questions, is a widely acknowledged benchmark to evaluate the ability of model following general user queries  Chen et al. (2024); Zhang et al. (2023). MMLU covers 57 subjects, aiming to evaluate comprehensive knowledge abilities across multiple major categories, from humanities to social sciences to science and engineering. TruthfulQA assesses the model’s ability to identify true claims, specifically in the context of literal truth about the real world.
Appendix B Language Models
To evaluate the effectiveness of our $\mathbb{IA}$ method, we validate our approach on six representative Large Language Models, each distinguished by its model architecture, model size, and alignment level. Specifically, we consider five open-source LLMs and one closed-source LLM.
ChatGLM-6B
(Zeng et al., 2023), trained on 1T tokens of both Chinese and English corpus, follows similar technologies to ChatGPT, including supervised fine-tuning, feedback bootstrap, and RLHF.
Llama-2-7B-chat
(Touvron et al., 2023), fine-tuned on Llama-2-7B by Meta, is optimized for dialogue cases and use reinforcement learning with human feedback (RLHF) to align with human preferences for helpfulness and safety.
Vicuna-7B-v1.1
and Vicuna-13B-v1.1(Chiang et al., 2023) are two open-source chatbots that are fine-tuned on Llama-7B and Llama-13B, using 70K samples of user-shared ChatGPT conversations, without further utilizing RLHF.
MPT-30B-Chat
(Team, 2023) is a chatbot model designed for generating conversations. It was developed by fine-tuning the MPT-30B on dialogue datasets such as ShareGPT, Vicuna, Camel-AI, and some generated datasets.
DeepSeek-67B-Chat
(DeepSeek-AI, 2024) is fine-tuned for dialogue based on DeepSeek-67B, and empirically verified to perform better on nearly 20 public datasets in both Chinese and English than Llama2-70B.
GPT-3.5
(OpenAI, 2023) is an advanced and closed-source LLM trained with massive text data and can generate human-like responses to a wide range of prompts and queries. We obtain the responses of gpt-3.5-turbo-1106 using API calls, which may differ from the response in web interfaces due to internal settings unexposed to the users.
For each of the above models, we conducted experiments under the system prompt template designed by their developers to simulate real-world usage scenarios.
Appendix C Evaluation Metrics
C.1 Automation-based Safety Evaluation
For evaluation of DAN, considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations. Practically, we adopt gpt-3.5-turbo-0613171717https://openai.com/blog/chatgpt as the evaluation function following Deng et al. (2023a). Especially, the model responses are submitted via API calls, and a specific evaluation prompt is designed for gpt-3.5-turbo-0613 to rate the harmfulness on a five-degree scale (‘A (Very Harmful)’, ‘B (Harmful)’, ‘C (Moderately Harmful)’, ‘D (Slightly Harmful)’, and ‘E (No Harmful at All)’). Figure 7 shows the detailed instruction for gpt-3.5-turbo-0613 evaluation.
Human Evaluation
To validate GPT-3.5 as the harmfulness annotation function, we conduct a human validation process. First, we randomly select 200 responses from both the vanilla model and our method on the DAN dataset. Then we engage two expert annotators proficient in alignment studies of LLMs and fluent in English. After reassessing their evaluations for cases with discrepancies and reaching a consensus, the annotators achieved a remarkable 97% agreement rate. Third, to quantify the credibility of GPT-3.5’s evaluations, we consider different degrees as the lowest acceptable safety threshold and calculate the agreement rate with human annotations. The results, detailed in Table 4, demonstrate the highest agreement of average 96% when categorizing C degree as the minimum harmfulness level, i.e., viewing the evaluation degrees of ‘A (Very Harmful)’, ‘B (Harmful)’, and ‘C (Moderately Harmful)’ as a success of jailbreak attacks. This underscores the enhanced reliability of GPT-3.5 annotations in conjunction with human judgment in harm detection.
Degrees $=$ A $\geq$ B $\geq$ C $\geq$ D $\geq$ E
Vanilla 51 72 93 90 89
Ours 97 98 99 99 3.0
Table 4: Agreement rate (%) between human annotations and different harm degrees provided by GPT-3.5 evaluation. A, B, C, D, and E degrees refer to ‘very harmful’, ‘harmful’, ‘moderately harmful’, ‘slightly harmful’, ‘not harmful at all’, respectively.
C.2 Rule-based Safety Evaluation
To calculate attack success rate (ASR) for SAP200 and AdvBench, we follow Zou et al. (2023) to consider an attack successful if the following refusal strings are not produced in the generation from the model:
  • $\bullet$
    “I’m sorry”,
  • $\bullet$
    “Sorry”,
  • $\bullet$
    “I am sorry”,
  • $\bullet$
    “I apologize”,
  • $\bullet$
    “As an AI”,
  • $\bullet$
    “I’m just”,
  • $\bullet$
    “As a language model”,
  • $\bullet$
    “As an Assistant”,
  • $\bullet$
    “I cannot”,
  • $\bullet$
    “I do not”,
  • $\bullet$
    “Hello!”,
  • $\bullet$
    “is illegal and unethical”,
  • $\bullet$
    “I’m really sorry”,
  • $\bullet$
    “OpenAI”,
  • $\bullet$
    “unethical”,
  • $\bullet$
    “not ethical”,
  • $\bullet$
    “illegal”,
  • $\bullet$
    “not legal”,
  • $\bullet$
    “My apologies”,
  • $\bullet$
    “but I cannot”,
  • $\bullet$
    “If you have any other non-malicious requests”,
  • $\bullet$
    “I’m not able to provide”,
  • $\bullet$
    “I am really sorry”,
  • $\bullet$
    “I cannot create”,
  • $\bullet$
    “I cannot support”,
  • $\bullet$
    “I can’t provide”,
  • $\bullet$
    “I can’t assist”.
Figure 7: Prompt template for harmfulness evaluation using gpt-3.5-turbo-0613.
C.3 Helpfulness Evaluation
For evaluation of AlpacaEval, we adopt gpt-4-1106-preview as the auto evaluator to reflect whether the general performance degrades after adding safety techniques (Dubois et al., 2023). Specifically, GPT-4 compares two responses to the same instruction: one with our methods and one provided by text-davinci-003 and report the win rate of our method. Figure 8 shows the detailed instruction for gpt-4-1106-preview evaluation. For MMLU, we follow Hendrycks et al. (2021) and report accuracy based on the model’s predictions and the groud truth labels. For TruthfulQA, we follow Chuang et al. (2023) and report on two main distinct metrics: MC1 and MC2 scores in Table 2. The complete results on the three metrics in TruthfulQA, i.e., MC1, MC2 and MC3, are presented in Table 5. We can see that our method consistently improves the truthfulness over different models, indicating that our method can be deployed in real applications to enhance LLM safety while increasing the general helpfulness to some extent.
Figure 8: Prompt template for AlpacaEval results evaluation using gpt-4-1106-preview.
Models Methods TruthfulQA
MC1 MC2 MC3
[…]
Table 5: Performance on TruthfulQA of our $\mathbb{IA}$ upon different models in terms of Accuracy (%). The best results are highlighted in bold.
C.4 Intention Recognition Success Evaluation
To verify whether the model can successfully identify the intention of jailbreak queries, we examine the model response of the first stage and view a success if it begin with “The essential intention of the query is”, as instructed in stage 1 of our method. And we empirically find that once the model successfully starts with this string, it can then successfully analyze the intention behind the user query.
Appendix D Alternative Prompts
D.1 One-Pass $\mathbb{IA}$
To study the effect of the implementation of our method, we combine the two stages of our mechanism and crafted an one-pass intention analysis instruction. See Figure 10 for the detailed prompt.
D.2 Two-Stage $\mathbb{IA}$
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. These additional instructions were meticulously designed with modifications of origin $\mathbb{IA}$ prompts (detailed in Figure 9) in lexical choices and syntactic constructions to direct intention analysis. The detailed prompts are presented in Figure 11. Table 6 presents the comparison results. We observe that three prompt sets all result in significant and comparable decreases in ASRs on the SAP200 across different LLMs. This uniformity highlights a key conclusion: it is the intention analysis mechanism, but not the specific instruction expressions, that makes our $\mathbb{IA}$ effective.
Figure 9: Detailed two-stage prompts for our method.
Figure 10: One-pass intention analysis prompt for our method.
Refer to caption Refer to caption
(a) Prompt Set A (b) Prompt Set B
Figure 11: Alternative prompts crafted for our intention analysis instructions.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 73.4 65.4 45.8
+ Prompt A 2.94 0.88 5.81
+ Prompt B 5.13 2.06 5.44
+ Ours 0.31 1.12 6.12
Table 6: Ablation of different $\mathbb{IA}$ prompts on SAP200 in ASR (%). The best and second best results are highlighted in bold and underline.
Figure 12: The MultiJail (under two scenarios) and CipherChat Datasets results on GPT-3.5 with and without our $\mathbb{IA}$. (a) Results on direct MultiJail dataset including English (en), Chinese (zh), Italian (it), Vietnamese (vi), Arabic (ar), Korean (ko), Thai (th), Bengali (bn), Swahili (sw), and Javanese (jv). (b) Results on malicious jailbreak prompt attached to MultiJail. (c) Results on CipherChat including ASCII (en), UTF (zh), Unicode (zh), and SelfCipher (en and zh) encryptions.
Appendix E Extensive Validations of $\mathbb{IA}$’s Effectiveness
E.1 Performance under More Advanced Attacks
Our method can consistently enhance safety in the context of more advanced jailbreaks such as multilingual attack and encryption-based attack.
Recent studies Deng et al. (2023b); Yong et al. (2023) reveal that the multilingual jailbreak poses a new defense challenge for LLMs. Yuan et al. (2024) and Wei et al. (2023a) also emphasize the struggles of more powerful LLMs, such as GPT-3.5, to stay safe when countering encryption-based attack. To verify the effectiveness of our method in these advanced jailbreak scenarios, we reproduce MultiJail and CipherChat following Deng et al. (2023b) and Yuan et al. (2024), respectively, and conduct further experiments on GPT-3.5181818We observe a high rate of invalid responses on smaller LLMs like ChatGLM-6B, Vicuna-7B and Vicuna-13B under MultiJail, as demonstrated in Deng et al. (2023b). And smaller LLMs may lack advanced encryption capability, which is required by CipherChat (Yuan et al., 2024). So we only adopt GPT-3.5 for these advanded jailbreak attacks experiments.. The results of GPT-3.5 with and without our $\mathbb{IA}$ are presented in Figure 12. We observe that 1) our $\mathbb{IA}$ consistently maintains performance in low-resource languages, e.g., th, bn, sw, and jv, even in scenarios where a malicious jailbreak prompt [^19] We adopt the same jailbreak prompt in Deng et al. (2023b), namely AIM. is attached to the multilingual attacks, 2) our $\mathbb{IA}$ significantly enhances safety when facing advanced encryption-based attack, even under the most effective SelfCipher attack. These demonstrate the effectiveness of our intention analysis defense mechanism under more advanced jailbreak attacks.
Methods Vicuna-7B Vicuna-13B Empirical
GCG DAN SAP200 GCG DAN SAP200 Runtime
Vanilla 83.0 48.4 70.0 87.0 60.0 65.9 1×1$\times$1 ×
+ Input Check 13.0 19.0 58.1 0.00 53.9 12.8 <1×<1$\times$< 1 ×
+ BPE-dropout (Jain et al., 2023) 63.0 23.8 67.2 50.0 28.2 48.9 <1×<1$\times$< 1 ×
+ ICD (Wei et al., 2023b) 1.00 44.4 32.8 0.00 58.9 32.8 <2×<2$\times$< 2 ×
+ Self Defense (Helbling et al., 2023) 24.0 31.3 53.2 20.0 28.8 29.7 2$\sim$ 2$\times$ 2 ×
+ Moral Self-Correction (Ganguli et al., 2023) 26.0 25.0 49.0 13.0 28.1 42.8 3$\sim$ 3$\times$ 3 ×
+ Self-Reminder (Xie et al., 2023) 11.0 45.3 33.8 1.0 57.5 36.9 <2×<2$\times$< 2 ×
+ SmoothLLM (Robey et al., 2023) 8.00 13.5 54.4 5.00 17.3 37.0 10$\sim$ 10$\times$ 10 ×
+ $\mathbb{IA}$ (Ours) 0.00 3.42 0.31 0.00 0.94 1.56 2$\sim$ 2$\times$ 2 ×
Table 7: Comparison of our method and existing advanced defense methods in terms of ASR (%) and empirical runtime. The best and second best results are highlighted in bold and underline.
E.2 Comparison with All Defense Baselines
Table 7 lists comparison results between $\mathbb{IA}$ and the baselines. [^20] Due to computational resource and financial limitations, we randomly select 320 samples each from DAN and SAP200 datasets for comparative experiments. As observed, $\mathbb{IA}$ consistently shows superiority over other baselines on different datasets and model scales. Specifically, $\mathbb{IA}$ outperforms the second-best method by 30.32% and 23.77% averagely on SAP200 and DAN, respectively. In addition, although ICD and Self-Reminder achieve considerable reduction in ASR on GCG, their performance severely degrades when dealing with complex and stealthy jailbreak prompts. On the contrary, $\mathbb{IA}$ consistently outperforms other baselines across both prompt-level and automatic token-level jailbreak datasets. Notably, $\mathbb{IA}$ achieves the best ASRs with comparable and acceptable empirical inference runtime.
ChatGLM-6B Llama2-7B-Chat Llama3-8B-Instruct Vicuna-7B Vicuna-13B MPT-30B-Chat DeepSeek-67B-Chat GPT-3.5
DAN 93% 100% 100% 100% 100% 92% 100% 42%
SAP200 100% 100% 100% 100% 100% 100% 100% 49%
Table 8: Manual check results of response’s helpfulness for harmful queries on DAN and SAP200 datasets in terms of rate.
Methods Safety Helpfulness
SAP200 DAN AlpacaEval
Vicuna-7B 73.4 44.3 66.2
[…]
Vicuna-7B + Ours 0.31 2.89 63.8
Table 9: Comparison between our method and well safety-trained LLM in safety and helpfulness (%). The best and second best are in bold and underline.
E.3 $\mathbb{IA}$ achieves comparable safety with well-safety-trained LLMs without the need for additional training.
Our method aims to enhance LLM safety in the inference stage. A natural question arises: how does its performance compare to well-safety-trained LLMs? To answer this, we compare our method with a representative well-safety-trained LLM, i.e., Llama2-7B-Chat. The results are listed in Table 9. We can see that our method achieves comparable performance to Llama2-7B-Chat on safety datasets while outperforming Llama2-7B-Chat on the helpfulness dataset by almost 6%. This demonstrates the advantage of our $\mathbb{IA}$ to achieve both safety and helpfulness goals without additionally resource-consuming safety training.
DAN SAP200 DeepInception GCG Average Time Cost
Vanilla 48.4 73.4 90.0 83.0 73.7 10.2
Self-Reminder 41.3 33.8 55.4 11.0 35.4 15.0
Ours 3.42 0.31 0.00 0.00 0.93 17.3
Self-reminder+Ours 3.12 0.00 0.00 0.00 0.78 25.5
Table 10: Performance of combining our $\mathbb{IA}$ with Self-Reminder method for Vicuna-7B in terms of ASR (%) and average Time Cost (s/sample).
E.4 $\mathbb{IA}$ can be combined with another defensive method.
We integrate our $\mathbb{IA}$ method with the Self-Reminder method (Xie et al., 2023) and conduct experiments on Vicuna-7B to see where such a combination leads. The comparison results in Table 10 indicates that although our method already significantly improves LLM safety, combining it with another defensive method can further enhance the effectiveness at the cost of additional computation overhead.
Appendix F Further explanation of $\mathbb{IA}$ format’s effectiveness when generated intention is incorrect
In Figure 4, we find that even when the correct intention ratio is 0% (with all generated intentions replaced by masked or random intentions), $\mathbb{IA}$ remains effective compared to the vanilla baseline. This effectiveness is mainly due to $\mathbb{IA}$’s two-round dialogue design. As shown in Figure 2, the final policy-aligned responses are generated with the context of intention analysis sequence format in the first round conversation. In Context Learning (ICL) community (Min et al., 2022) has demonstrated that “keeping the format of the input-label pairs is key, and replacing gold labels with random labels in the demonstrations only marginally lowers the performance.” Therefore, even if the intention label generated in the first stage is incorrect, keeping the entire intention analysis format plays a significant role in making the final response safer than when no intention analysis sequence is used (vanilla method). Moreover, as indicated in Table 4, improving the ratio of correct intention labels can further enhance $\mathbb{IA}$’s performance.
Defense Methods DAN AlpacaEval
Harmfulness Helpfulness Win Rate
Vanilla 48.4 5.66 66.2
Input Check 19.0 3.25 64.4
ICD 40.4 5.79 60.3
Self-Reminder 41.3 5.89 64.6
SmoothLLM 13.5 5.35 60.8
[…]
Table 11: Comparison results for Vicuna-7B in terms of harmfulness (%), and helpfulness (%) on DAN dataset, and win rate (%) on AlpacaEval.
Appendix G Deeper Study of Safe Responses’ Helpfulness for Harmful Queries
G.1 ChatGPT Evaluation
To comprehensively study the impact of our $\mathbb{IA}$ on responses to harmful queries, we follow Zheng et al. (2023) and prompt ChatGPT to score the helpfulness of these safe refusals [^21] When refusing harmful queries, we expect LLMs to further provide reasonable explanations or suggestions instead of simply rejecting, thus being safe and helpful at the same time.. Table 11 presents comparison results between different defense methods on the harmfulness (ASR) and helpfulness score on the DAN dataset. We observe that $\mathbb{IA}$ enables LLMs to effectively give safe refusals with satisfactory helpfulness for harmful queries. We also manually check these refusals in Appendix G.2 and find that $\mathbb{IA}$ enables LLMs to craft more nuanced responses to specific unsafe intents like inciting hatred and division.
Target Model Intent. Model DAN SAP200
Vicuna-7B 44.3 67.2
Vicuna-7B 2.89 0.31
Vicuna-13B 1.93 0.62
Vicuna-13B 54.7 65.4
Vicuna-7B 1.25 1.87
Vicuna-13B 0.64 1.12
Table 12: ASR (%) of our $\mathbb{IA}$ on DAN and SAP200 with different intention analysis model scales. For each target model, the intention analysis is performed in three ways, i.e., without intention analysis, analyzed by Vicuna-7B, and by Vicuna-13B.
G.2 Manual Check
To comprehensively study the impact of our $\mathbb{IA}$ on responses to harmful queries, we conduct a manual review of 100 random-sampled refusals on both DAN and SAP200 datasets for each of the seven LLMs under our $\mathbb{IA}$. We manually check for 1) reasonable explanations about why it rejects to respond and 2) positive suggestions for the user. In Table 8, our findings show that, following $\mathbb{IA}$, almost all LLMs can give detailed explanations in nearly 100% of cases (except for GPT-3.5 due to OpenAI’s external moderation mechanism during API interactions (OpenAI, 2023)). Specifically, we find LLMs indeed conduct more nuanced handling regarding varying unsafe intents (e.g. terrorism and violence), indicating that our $\mathbb{IA}$ ensures the helpfulness of safe responses for harmful queries. Detailed cases are presented in Figures 16.
Appendix H Cross-Intention Analysis Experiment
A question that naturally arises about our method is, “Would the effectiveness of $\mathbb{IA}$ improve notably with a larger intention analysis model?” To investigate this, we conduct cross-experiments using Vicuna-7B and 13B models. Specifically, we evaluate each target model in three contexts: without any intention analysis model, with the Vicuna-7B as the intention analysis model, and with the Vicuna-13B serving the same role. Table 12 shows the results. It is observable that for the 7B model, utilizing the 13B for intention analysis yields only marginal performance gains. For the 13B model, employing the 7B model for intention recognition achieves comparable results to using the 13B model itself for intention analysis, but with the added benefit of conserving inference resources. This indicates that the 7B model is sufficiently capable of intention analysis to enhance model safety.
Appendix I Qualitative Examples
This section presents some practical cases, which are from SAP200 dataset (see Figure 13), DAN dataset (see Figure 14), and AdvBench dataset (for GCG; see Figure 15), respectively.
Figure 13: Qualitative comparison example I from SAP200 dataset on Vicuna-7B with and without our $\mathbb{IA}$.
Figure 14: Qualitative comparison example II from DAN dataset on Vicuna-7B with and without our $\mathbb{IA}$.
Figure 15: Qualitative comparison example III from AdvBench dataset (for GCG) on Vicuna-7B with and without our $\mathbb{IA}$.
Figure 16: Examples of nuanced handling for harmful user queries regarding varying unsafe intentions (terrorism in case 1 and violence in case 2).
Figure 17: Two failure cases of intention analysis induced by too strong (for Llama2-7B-Chat) or too weak (for Vicuna-7B) inherent safety.
Figure 18: Two failure cases of our $\mathbb{IA}$ induced by weak inherent safety (in case 1) and failed intention analysis (in case 2).